OpenStudy (anonymous):

Calculate the derivative of the function. Then find the value of the derivative as specified. f(x) =8/x ; f'(-1)

OpenStudy (anonymous):

\[f(x)= 8/x = 8*x ^{-1}\] differentiate yields: \[f \prime(x)= (-1)*8*x ^{-2} = -8/x ^{2}\] to find f'(-1), just substitute x=-1 in the derivative